OpenStudy (anonymous):

What is the vertex of the parabola whose equation is y = (x + 1)2 + 3?

jigglypuff314 (jigglypuff314):

\[\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\ \Huge\sf\color{blue}{~~~~Welcome ~To~ OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color {blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\] vertex form for a parabola: y = (x - h)^2 + k when coordinate (h, k) is the vertex so for y = (x + 2)^2 + 3 ^ ^ so h = -2 and k = 3